When someone is admitted to the hospital for severe dehydration, one treatment is to start an I.V. of Hartmann's solution to replace body fluids and salts.

Nurses often prepare these IVs from concentrated solutions. The total desired ion concentration for the I.V. is 276 mM. For a 500mL I.V. of Hartmann's solution, what volume of the 1.38M stock is needed?

100. mL
200. mL
500. mL
50.0 mL
10.0 mL
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 1.38 M

M2 = 0.276 M

V2 = 500 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.276*500)/1.38

V1 = 100 mL

Answer: 100 mL